OpenStudy (anonymous):

expand and simplify 9a+3(8-2a)

OpenStudy (anonymous):

Good afternoon, my beloved child.

OpenStudy (anonymous):

Do you know the distributive property?

OpenStudy (anonymous):

\( \color{Black}{\Rightarrow a(b \pm c) = ab \pm ac }\)

OpenStudy (anonymous):

whats that?

OpenStudy (anonymous):

It is a property that is applied on all numbers, including algebraic ones.

OpenStudy (anonymous):

\(3(8 - 2a) = (3\times 8) - (3 \times 2a)\)

OpenStudy (anonymous):

nope i have no idea, is there any chance you could please teach me how to expand and simplify and is there an easy way of remembering it?

OpenStudy (anonymous):

3(a+8) is the answer .

OpenStudy (anonymous):

If you work at a restaurant, and you have water in a jug, you distribute the water to all the glasses. Similar with numbers.

OpenStudy (anonymous):

\( \color{Black}{\Rightarrow3(8 - 2a) = (3 \times 8) - (3 \times 2a) = 24 - 6a }\)

OpenStudy (anonymous):

This will simplify to \(9a + 24 - 6a\) Combine the like terms. \( \color{Black}{\Rightarrow 3a + 24 }\) If you need to factor, then you may. \( \color{Black}{\Rightarrow 3(a + 8) }\)

OpenStudy (anonymous):

way too cumbersome isaac .

OpenStudy (anonymous):

okay i am starting to get it

OpenStudy (anonymous):

@kimvallance Keep it up, my child.

OpenStudy (anonymous):